OpenStudy (anonymous):

For the equation 2x 3 - 11x 2 + 12x + 9 = 0, the root 3 has a multiplicity of _____.

OpenStudy (anonymous):

\[2x^{3}-11x^{2}+12x+9 =(x-3)^{2}(2x+1)=0\]Now multiplicity means the exponent of the product, Thus what is the exponent of (x-3)?

OpenStudy (anonymous):

x2-6x+9=0